OpenStudy (anonymous):

WHAT is (5.3/100)(3)

OpenStudy (mew55):

0.159

OpenStudy (anonymous):

yeah is 0.159

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{5.3}{100}\right)^3\implies \cfrac{5.3^3}{100^3}\)

OpenStudy (anonymous):

I am doing a problem that is B=P+Prt

OpenStudy (mew55):

divide 5.3/100 first =0.053 then multiply by 3

OpenStudy (anonymous):

and i replace the variables with the number so i got 6954=P+P(5.3/100)(3)

OpenStudy (anonymous):

So then i got 6954=P+(0.053) I subtracted 0.053 on both sides and got

OpenStudy (anonymous):

Sorry it was 0.159

OpenStudy (anonymous):

and i got a huge number. which i know its wrong

OpenStudy (anonymous):

what did i do wrong ?

OpenStudy (mew55):

where the 6954 come from?

OpenStudy (anonymous):

its a problem solving. Its says. Suppose that an account earns 5.3% interest and after 3 years the balance is 6954$.Find the principle that was invested.Use the formula B=P+Prt,which is used to calculate the balance in an account if simple interest is added to a principle P at an interest rate of r after t years.

OpenStudy (anonymous):

@cupcakejin please help me